OpenStudy (anonymous):

Given that the density of TlCl(s) is 7.00 g/cm3 and that the length of an edge of a unit cell is 385 pm, determine how many formula units of TlCl there are in a unit cell.

OpenStudy (anonymous):

Density= ZM/a^3 N. Here Z is the number of effective Ti(formula units). a=edge length. N=avogadro number=6.022*10^23. M=Molecular mass. You know everything excpet Z. So you can substitute and find it.

OpenStudy (anonymous):

I think it is 5. but check the calculations.